### v3

# PanIC: consistent information criteria for general model selection problems

Abstract
Model selection is a ubiquitous problem that arises in the application of many statistical and machine learning methods. In the likelihood and related settings, it is typical to use the method of information criteria (IC) to choose the most parsimonious among competing models by penalizing the likelihood-based objective function. Theorems guaranteeing the consistency of IC can often be difficult to verify and are often specific and bespoke. We present a set of results that guarantee consistency for a class of IC, which we call PanIC (from the Greek root ’pan’, meaning ’of everything’), with easily verifiable regularity conditions. The PanIC are applicable in any loss-based learning problem and are not exclusive to likelihood problems. We illustrate the verification of regularity conditions for model selection problems regarding finite mixture models, least absolute deviation and support vector regression, and principal component analysis, and we demonstrate the effectiveness of the PanIC for such problems via numerical simulations. Furthermore, we present new sufficient conditions for the consistency of BIC-like estimators and provide comparisons of the BIC to PanIC.
School of Computing, Engineering and Mathematical Sciences, La Trobe University, Bundoora, Australia. Institute of Mathematics for Industry, Kyushu University, Fukuoka, Japan.
Email: h.nguyen5@latrobe.edu.au.
Keywords: Information criteria; model selection; order selection; loss minimization; finite mixture models; least absolute deviation; support vector regression; principal component analysis.
1 Introduction
Let $\left(\Omega,\mathcal{F},\text{Pr}\right)$ be a probability space with typical element $\omega$ and expectation operator E, and suppose that $X:\Omega\rightarrow\mathbb{X}\subset\mathbb{R}^{d}$ is a random variable on the image probability space $\left(\mathbb{X},\mathcal{B}\left(\mathbb{X}\right),\text{Pr}_{X}\right)$, with expectation operator $\text{E}_{X}$, where $\mathcal{B}\left(\mathbb{X}\right)$ is the Borel $\sigma$-algebra of $\mathbb{X}$. Suppose that we observe an independent and identically distributed (IID) sequence $\left(X_{i}\right)_{i\in\left[n\right]}$, where $X_{i}$ has measure $\text{Pr}_{X}$, for each $i\in\left[n\right]=\left\{1,\dots,n\right\}$ ($n\in\mathbb{N}$).
Let $\left(\mathscr{H}_{k}\right)_{k\in\left[m\right]}$ be a sequence of hypotheses that define functional spaces of models determined by a parameter vector $\theta_{k}\in\mathbb{T}_{k}\subset\mathbb{R}^{q_{k}}$, where $q_{k}\in\mathbb{N}$, for each $k\in\left[m\right]$. That is, we can write:
$\mathscr{H}_{k}=\left\{h_{k}\left(\cdot;\theta_{k}\right):\mathbb{X}\rightarrow\mathbb{R}\text{, }\theta_{k}\in\mathbb{T}_{k}\right\}\text{.}$
Further, define a loss function $\ell:\mathbb{R}\rightarrow\mathbb{R}$ and, for each $k\in\left[m\right]$, call
$R_{k,n}\left(\theta_{k}\right)=\frac{1}{n}\sum_{i=1}^{n}\ell\left(h_{k}\left(X_{i};\theta_{k}\right)\right)$
and
$r_{k}\left(\theta_{k}\right)=\text{E}_{X}\left\{\ell\left(h_{k}\left(X;\theta_{k}\right)\right)\right\}$
the empirical and expected risks, defined by parameter $\theta_{k}$, within hypothesis $\mathscr{H}_{k}$, respectively. For brevity, we will write $\ell_{k}\left(X;\theta_{k}\right)$ in place of $\ell\left(h_{k}\left(X;\theta_{k}\right)\right)$.
Define
$k^{*}=\min\underset{k\in\left[m\right]}{\arg\min}\left\{\min_{\theta_{k}\in\mathbb{T}_{k}}r_{k}\left(\theta_{k}\right)\right\}$
to be the optimal hypothesis index. Here, it is assumed that there is an order of ‘complexity’ that is indexed by $k$, such as when $\left(\mathscr{H}_{k}\right)_{k\in\left[m\right]}$ is a nested sequence (i.e., $\mathscr{H}_{k}\subset\mathscr{H}_{k+1}$, for each $k\in\left[m-1\right]$), or when the parameter spaces $\left(\mathbb{T}_{k}\right)_{k\in\left[m\right]}$ have increasing dimensions (i.e., $q_{k}<q_{k+1}$, for each $k\in\left[m-1\right]$), which may correspond to both nested and non-nested sequences of hypotheses. Our objective is to obtain a sequence of estimates $\left(\hat{K}_{n}\right)_{n\in\mathbb{N}}$ of $k^{*}$, where $\hat{K}_{n}:\mathbb{X}^{n}\rightarrow\left[m\right]$, for each $n\in\mathbb{N}$, and where $\hat{K}_{n}$ has the consistency property:
$\lim_{n\rightarrow\infty}\text{Pr}\left(\hat{K}_{n}=k^{*}\right)=1\text{.}$ (1)
The problem of estimating $k^{*}$ has long been studied in statistics and machine learning under the topics of model identification, model selection, variable selection, and structural risk minimization, among other names, where $\mathscr{H}_{k^{*}}$ is often referred to as the class of parsimonious models. We will generically refer to the problem as model selection, without loss of generality. For treatments on the subject, we refer the interested reader to the volumes of McQuarrie and Tsai (1998), Burnham and Anderson (2002), Massart (2007), Claeskens and Hjort (2008), Konishi and Kitagawa (2008), Ando (2010), Gassiat (2018), and Oneto (2020), as well as the expositions of Vapnik (1998, Ch. 6), Leeb and Potscher (2005), Leeb and Potscher (2009), Mohri et al. (2018, Ch. 4), and Hansen (2022, Ch. 28).
In this work, we consider the method of penalization, best identified with information criteria (IC), that was pioneered in the works of Akaike (1974) and Schwarz (1978). That is, we propose estimators of $k^{*}$ of the form
$\hat{K}_{n}=\min\underset{k\in\left[m\right]}{\arg\min}\text{ }\left\{\min_{\theta_{k}\in\mathbb{T}_{k}}R_{k,n}\left(\theta_{k}\right)+P_{k,n}\right\}\text{,}$ (2)
where $P_{k,n}:\Omega\rightarrow\mathbb{R}_{\geq 0}$ are random variables dependent on the data and hypothesis index, which we shall refer to as penalties.
The reader will be able to identify this framework with the Akaike IC (AIC; Akaike, 1974) and the Bayesian IC (BIC; Schwarz, 1978), respectively, by setting $\ell$ to be the negative logarithm, $\left(\mathscr{H}_{k}\right)_{k\in\left[m\right]}$ to be sets of probability density functions, and by taking $P_{k,n}$ to be equal to either $n^{-1}q_{k}$ or $n^{-1}q_{k}\log n/2$.
Let $\theta_{k}^{*}$ be a minimizer of $r_{k}$, for each hypothesis $\mathscr{H}_{k}$. In Sin and White (1996), general consistency results (i.e., sufficient conditions guaranteeing (1)) are provided for broad classes of losses and penalties, under strong assumptions regarding $\theta_{k}^{*}$, $\ell$, and $h_{k}$. For example, $\theta_{k}^{*}$ is required to be unique and an interior point of $\mathbb{T}_{k}$, for each $n$, and $\ell_{k}\left(x;\cdot\right)$ must be twice continuously differentiable, for each $x\in\mathbb{X}$, where $\partial^{2}r\left(\theta\right)/\partial\theta\partial\theta^{\top}$ is assumed to exist and be positive definite at $\theta_{k}^{*}$. We note that comparable general results were obtained in Baudry (2015).
We generalize the results of Sin and White (1996) and Baudry (2015) by providing conditions under which we may admit non-uniqueness of $\theta_{k}^{*}$ and lack of differentiability of $\ell_{k}\left(x;\cdot\right)$. This includes not requiring the positive-definiteness assumption on $\partial^{2}r\left(\theta_{k}^{*}\right)/\partial\theta\partial\theta^{\top}$, stated above. This permits simple constructions of information criteria for problems where such assumptions are non-trivial, such as for finite mixture models, as we discuss in the sequel. Other models where such pathologies occur are the so-called singular models discussed in Watanabe (2009) and Drton and Plummer (2017), which include common models such as factor analyzers, Bayesian networks and neural networks with hidden nodes, reduced rank regression models, and variants of such models, among others. This is achieved via asymptotic results for sample average approximations of stochastic programming problems, as covered in Shapiro (2000) and Shapiro et al. (2021, Ch. 5). Due to the generality of our result for consistent estimation of $k^{*}$ across different modeling and learning problems, and due to our use of the IC form (2), we name our approach PanIC, using the Greek prefix ‘pan’ to mean ‘all’ or ‘of everything’.
Consistency results for IC have been studied in specifics in numerous previous works. For example, one may point to the thorough analyses of Leroux (1992), Keribin (2000), Hui et al. (2015), Drton and Plummer (2017), and Gassiat (2018, Ch. 4), when $\ell$ is taken as the negative logarithm and $\mathscr{H}_{k}$ are classes of finite mixture models; Varin and Vidoni (2005), Gao and Song (2010), Ng and Joe (2014), and Hui (2021), when $\ell_{k}$ are taken to be the negative composite, pseudo-, or quasi-log-likelihood objects; and Kobayashi and Komaki (2006), Claeskens et al. (2008), and Zhang et al. (2016), in the support vector machine context. Furthermore, the finite sample properties of the IC has been studied in many situations, for example, in the works of Barron et al. (1999), Massart (2007), Bartlett (2008), and Giraud (2022). Instead of seeking to replace the thorough and specific treatments provided by the listed texts, we aim to present general results for obtaining consistent estimators of $k^{*}$ in situations where other approaches may be difficult to verify or are unavailable.
The rest of the work proceeds as follows. We provide an exposition of the PanIC approach and its theoretical properties in Section 2. Example applications are provided in Section 3 and numerical experiments are performed in Section 4. We discuss our work in Section 5. Proofs and technical results are relegated to the Appendix of the text.
2 Main results
We shall retain the notation and technical setting from Section 1 and make the following assumptions, for each $k\in\left[m\right]$:
A1
$\mathbb{T}_{k}$ is compact and there exists a $\tau_{k}\in\mathbb{T}_{k}$, such that
$\text{E}\left\{\ell_{k}\left(X;\tau_{k}\right)^{2}\right\}<\infty\text{.}$
A2
There exists a measurable function $\mathfrak{C}_{k}:\mathbb{X}\rightarrow\mathbb{R}_{\geq 0}$, such that $\text{E}\left\{\mathfrak{C}_{k}\left(X\right)^{2}\right\}<\infty$ and
$\left|\ell_{k}\left(x;\theta_{k}\right)-\ell_{k}\left(x;\tau_{k}\right)\right|\leq\mathfrak{C}_{k}\left(x\right)\left\|\theta_{k}-\tau_{k}\right\|\text{,}$
for every $\theta_{k},\tau_{k}\in\mathbb{T}_{k}$ and almost every $x\in\mathbb{X}$.
Notice that A1 and A2 together imply that $\ell_{k}\left(x;\theta_{k}\right)$ is Caratheodory in the sense that $\ell_{k}\left(x;\cdot\right):\mathbb{T}_{k}\rightarrow\mathbb{R}$ is continuous for each $x\in\mathbb{X}$, and $\ell_{k}\left(\cdot;\theta_{k}\right):\mathbb{X}\rightarrow\mathbb{R}$ is measurable for each $\theta_{k}\in\mathbb{T}_{k}$, for each $k\in\left[m\right]$. We write
$\mathbb{T}_{k}^{*}=\arg\min_{\theta_{k}\in\mathbb{T}_{k}}\text{ }r_{k}\left(\theta_{k}\right)$
to denote the set of minima of the expected risk under hypothesis $\mathscr{H}_{k}$. We shall indicate convergence in distribution by the symbol $\rightsquigarrow$. A direct application of Shapiro et al. (2021, Thm. 5.7) yields the following lemma, which is instrumental in the proof of the main result in the following section.
Lemma 1.
Assume that $\left(X_{i}\right)_{i\in\left[n\right]}$ is an IID sequence, and that A1 and A2 hold, for each $k\in\left[m\right]$. Then, for each $k$,
$\sqrt{n}\left(\min_{\theta_{k}\in\mathbb{T}_{k}}R_{k,n}\left(\theta_{k}\right)-\min_{\theta_{k}\in\mathbb{T}_{k}}r_{k}\left(\theta_{k}\right)\right)\rightsquigarrow\inf_{\theta_{k}\in\mathbb{T}_{k}^{*}}Y_{k}\left(\cdot;\theta_{k}\right)\text{,}$ (3)
where $Y_{k}$ is a bounded random process, indexed by $\mathbb{T}_{k}$, with continuous sample paths $Y_{k}\left(\omega;\cdot\right)$.
Remark 1.
We note that details regarding $Y_{k}$ are immaterial, since $Y_{k}\left(\omega;\theta_{k}\right)$ must be a bounded and continuous function of $\theta_{k}\in\mathbb{T}_{k}$, and thus $\inf_{\theta_{k}\in\mathbb{T}_{k}^{*}}Y_{k}\left(\omega;\theta_{k}\right)$ is finite for every $\omega\in\Omega$. Therefore $\inf_{\theta_{k}\in\mathbb{T}_{k}^{*}}Y_{k}\left(\cdot;\theta_{k}\right):\Omega\rightarrow\mathbb{R}$ is a random variable in the usual sense. Since the left-hand side (LHS) of (3) converges in distribution to a random variable, this permits us to conclude that the LHS of (3) is bounded in probability, as $n\rightarrow\infty$ (cf. Boos and Stefanski 2013, Sec. 5.5.3c).
2.1 PanIC
Make the following additional assumptions, for each $k\in\left[m\right]$:
B1
$P_{k,n}>0$, for each $k\in\left[m\right]$ and $n\in\mathbb{N}$, and $P_{k,n}=o_{\text{Pr}}\left(1\right)$, as $n\rightarrow\infty$.
B2
If $k<l$, then $\sqrt{n}\left\{P_{l,n}-P_{k,n}\right\}\rightarrow\infty$, in probability, as $n\rightarrow\infty$.
We will say that any estimator of form (2), satisfying B1 and B2, is a PanIC. The proof of the following result appears in Appendix B, along with those of Theorems 2 and 3.
Theorem 1.
Assume that $\left(X_{i}\right)_{i\in\left[n\right]}$ is an IID sequence, and that A1, A2, B1 and B2 hold, for each $k\in\left[m\right]$. Then, the PanIC (2) satisfies the consistency property (1).
Remark 2.
We note that PanIC were considered in Sin and White (1996, Prop. 4.6(a,b)). However, as noted in Section 1, the Sin and White (1996) results required much stronger assumptions than A1 and A2. Namely, it is assumed that both $\ell_{k}\left(x;\cdot\right)$ and $r_{k}$ are both twice continuously differentiable on a set $\mathbb{T}_{k}$, with non-empty interior, for each $k\in\left[m\right]$; that $r_{k}$ has a unique minimizer at $\theta_{k}^{*}$, in the interior of $\mathbb{T}_{k}$, for each $n\in\mathbb{N}$; that $\ell_{k}\left(X;\cdot\right)$ and its derivative $\partial\ell_{k}\left(X;\theta_{k}\right)/\partial\theta_{k}$ both satisfy central limit theorems; and the Hessian $\partial^{2}r_{k}\left(\theta_{k}\right)/\partial\theta_{k}\partial\theta_{k}^{\top}$ is positive definite at $\theta_{k}^{*}$. In contrast, Theorem 1 makes no assumptions regarding the uniqueness or even countability of minima of $r_{k}$ and $R_{k,n}$, nor the location of the minima within $\mathbb{T}_{k}$. Moreover, the totality of A1 and A2 amount to a Lipschitz assumption on $\ell_{k}\left(x;\cdot\right)$, and thus demands less than one derivative on each of the $k$ functions, although if $\ell_{k}\left(x;\cdot\right)$ are each continuously differentiable then the verification of A1 and A2 becomes procedural, as one can take the supremum norm of the gradients, with respect to $\theta_{k}$, in place of $\mathfrak{C}_{k}\left(x\right)$.
Remark 3.
Let $\left(c_{k}\right)_{k\in\left[m\right]}$ be positive constants, such that $c_{k}<c_{l}$, for each $k<l$, and define $\log_{+}\left(\cdot\right)=\max\left\{1,\log\left(\cdot\right)\right\}$. We are particularly interested in penalties taking the form:
$P_{n,k}=\alpha c_{k}\sqrt{\frac{\log_{+}^{\left(\beta\right)}\left(n\right)}{n}}\text{,}$ (4)
where $\alpha>0$, $\beta\in\mathbb{N}$, and
$\log_{+}^{\left(\beta\right)}\left(n\right)=\underset{\beta\text{ times}}{\underbrace{\left(\log_{+}\circ\log_{+}\circ\dots\circ\log_{+}\right)}}\left(n\right)\text{.}$
We call any IC with penalty of form (4) a SWIC (Sin–White IC) of order $\beta$, with constant $\alpha$, in homage to Sin and White (1996), who suggest the use of the first and second order versions.
We can easily check that penalties of form (4) satisfy B1 and B2, and thus SWIC are within the class of PanIC. B1 is verified by noting that $\log_{+}^{\left(\beta\right)}\left(n\right)$ is increasing and real-valued and increases slower than $\sqrt{n}$. B2 can be checked by observing that
$\sqrt{n}\left\{P_{l,n}-P_{k,n}\right\}=\alpha\left(c_{l}-c_{k}\right)\sqrt{\log_{+}^{\left(\beta\right)}\left(n\right)}\underset{n\rightarrow\infty}{\longrightarrow}\infty\text{,}$
since $\alpha\left(c_{l}-c_{k}\right)>0$, for $k<l$, as required.
2.2 BIC-like criteria
We shall say that a criterion is BIC-like if it satisfies the following assumption, along with B1, for each $k\in\left[m\right]$:
B2
If $k<l$, then $n\left\{P_{l,n}-P_{k,n}\right\}\rightarrow\infty$, in probability, as $n\rightarrow\infty$.
General conditions under which BIC-like estimators (2) are consistent were also studied in Sin and White (1996, Prop. 4.2(c)) and Baudry (2015, Thm. 8.1 and Cor. 8.2). Using further results from the stochastic programming literature, we can obtain alternative consistency criteria that are not implied by previous works. We make the following additional assumptions, for each $k\in\left[m\right]$:
C1
$r_{k}$ is Lipschitz continuous on $\mathbb{T}_{k}$, and is uniquely minimized and twice differentiable at some $\theta_{k}^{*}\in\mathbb{T}_{k}$.
C2
$\ell_{k}\left(x;\cdot\right)$ is Lipschitz continuous and differentiable at $\theta_{k}^{*}$, for almost all $x\in\mathbb{X}$.
C3
The set $\mathbb{T}_{k}$ is second order regular at $\theta_{k}^{*}$ (see Appendix B).
C4
The quadratic growth condition holds at $\theta_{k}^{*}$ (see Appendix B).
C5
If $r\left(\theta_{k}^{*}\right)=r\left(\theta_{k^{*}}^{*}\right)$, then $n\left(R_{k,n}\left(\theta_{k}^{*}\right)-R_{k^{*},n}\left(\theta_{k^{*}}^{*}\right)\right)=O_{\text{Pr}}\left(1\right)$.
Theorem 2.
Assume that $\left(X_{i}\right)_{i\in\left[n\right]}$ is an IID sequence, and that A1, A2, B1, B2, and C1–C5 hold, for each $k\in\left[m\right]$. Then, the BIC-like estimator (2) satisfies the consistency property (1).
Remark 4.
Sin and White (1996, Prop. 4.2(c)) prove the consistency of BIC-like criteria under the same assumptions as those discussed in Remark 2. A slight relaxation is provided in Baudry (2015, Thm. 8.1 and Cor. 8.2), where each $\ell_{k}\left(x;\cdot\right)$ is only assumed to be continuously differentiable on $\mathbb{T}_{k}$, for almost all $x\in\mathbb{X}$, for $k\in\left[m\right]$. However, it is still assumed that $\theta_{k}^{*}$ is in the interior of $\mathbb{T}_{k}$, and that the Hessian $\partial^{2}r_{k}\left(\theta_{k}\right)/\partial\theta_{k}\partial\theta_{k}^{\top}$ is positive definite at $\theta_{k}^{*}$, for each $k$. From Appendix A, we observe that Theorem 2 makes the same conclusions as the result of Baudry (2015), under the same assumptions on $\ell_{k}$ and $r_{k}$. However Theorem 2 provides further relaxations, allowing $\ell\left(x;\cdot\right)$ to be Lipschitz on $\mathbb{T}_{k}\backslash\left\{\theta_{k}^{*}\right\}$, and only requiring differentiability at $\theta_{k}^{*}$, for almost all $x\in\mathbb{X}$. Furthermore, $\theta_{k}^{*}$ is permitted to be on the boundary of $\mathbb{T}_{k}$, with C3 and C4 being particularly easy to verify when $\mathbb{T}_{k}$ are polyhedral sets. We note that the result of Sin and White (1996) is proved by first verifying that $\theta_{k}^{*}$ is asymptotically normal for each $k\in\left[m\right]$. An asymptotic expansion is then used to obtain the boundedness in probability of the $n$-scaled minimum risk. Alternatively, Baudry (2015) takes an empirical processes approach, where concentration inequalities for relevant terms are obtained and used to obtain the boundedness in probability of required terms.
Remark 5.
In Baudry (2015, Cor. 8.2), it is proposed that C5 can be implied by the more explicit assumption that $r_{k}\left(\theta_{k}^{*}\right)=r_{k^{*}}\left(\theta_{k^{*}}^{*}\right)$ if and only if $\ell_{k}\left(x;\theta_{k}\right)=\ell_{k^{*}}\left(x;\theta_{k^{*}}^{*}\right)$, for almost all $x\in\mathbb{X}$. Sin and White (1996) do not propose any method for checking C5 but suggests that it is generally satisfied in situations where the hypotheses are nested, in the sense that $\mathscr{H}_{k}\subset\mathscr{H}_{l}$, for each $k<l$, particularly when $\ell$ is taken to be the negative log-likelihood and $\mathscr{H}_{k}$ are sets of probability density functions, for each $k\in\left[m\right]$ (cf. Vuong, 1989).
Remark 6.
Of course our reference to ICs with penalties satisfying B1 and B2 as BIC-like is suggestive that the BIC should fall within this category. Upon assuming that $q_{k}<q_{l}$ for each $k<l$, the original BIC penalty of Schwarz (1978) corresponds to the choice $P_{k,n}=q_{k}n^{-1}\log n/2$, which verifies B2 by observing that
$n\left\{P_{l,n}-P_{k,n}\right\}=\left(q_{l}-q_{k}\right)\log n/2\underset{n\rightarrow\infty}{\longrightarrow}\infty\text{,}$
since $q_{l}-q_{k}>0$. Note that the class of BIC-like criteria is not limited to the BIC and include, for example, the Hannan–Quinn information criterion with penalty $P_{k,n}=q_{k}n^{-1}\log_{+}^{\left(2\right)}\left(n\right)/2$ (Hannan and Quinn, 1979), or more generally, IC with penalties of the form $P_{k,n}=\alpha q_{k}n^{-1}\log_{+}^{\left(\beta\right)}\left(n\right)$, for $\alpha>0$ and $\beta\in\mathbb{N}$.
3 Example applications
3.1 Finite mixture models
For each $k\in\left[m\right]$, let
$\mathscr{H}_{k}=\left\{\begin{array}[]{c}h_{k}\left(\cdot;\theta_{k}\right)=\sum_{z=1}^{k}\pi_{z}f\left(\cdot;\upsilon_{z}\right):\theta_{k}=\left(\pi_{1},\dots,\pi_{k},\upsilon_{1},\dots,\upsilon_{k}\right)\text{,}\\ \theta_{k}\in\mathbb{T}_{k}=\mathbb{S}_{k-1}\times\mathbb{U}^{k}\end{array}\right\}\text{,}$ (5)
where $f\left(\cdot;\upsilon\right)$ is a (component) density with parameter $\upsilon\in\mathbb{U}\subset\mathbb{R}^{q}$, where $\mathbb{U}$ is compact and $\mathbb{S}_{k-1}$ is the probability simplex in $\mathbb{R}^{k}$. We call $\mathscr{H}_{k}$ the set of finite mixtures of component densities $f\left(x;\upsilon\right)$ of order $k$.
Suppose that we observe IID data $\left(X_{i}\right)_{i\in\left[n\right]}$ from the density
$h_{k^{*}}\left(x;\theta_{k^{*}}^{*}\right)=\sum_{z=1}^{k^{*}}\pi_{z}^{*}f\left(x;\upsilon_{z}^{*}\right)\in\mathscr{H}_{k^{*}}\text{,}$
for some $k^{*}\in\left[m\right]$, where $\theta_{k}^{*}=\left(\pi_{1}^{*},\dots,\pi_{k}^{*},\upsilon_{1}^{*},\dots,\upsilon_{k}^{*}\right)\in\mathbb{T}_{k}$, for each $k\in\left[m\right]$, and
$\theta_{k}^{*}\in\underset{\theta_{k}\in\mathbb{T}_{k}}{\arg\min}\text{ }\underset{=r_{k}\left(\theta_{k}\right)}{\underbrace{\text{E}_{X}\left\{-\log h_{k}\left(X;\theta_{k}\right)\right\}}}\text{.}$ (6)
In the classical IC setting, if $k^{*}$ is unknown, then we typically estimate it via a penalized estimator of the form
$\hat{K}_{n}=\min\underset{k\in\left[m\right]}{\arg\min}\text{ }\left\{\min_{\theta_{k}\in\mathbb{T}_{k}}\underset{=R_{k,n}\left(\theta_{k}\right)}{\underbrace{-\frac{1}{n}\sum_{i=1}^{n}\log h_{k}\left(X_{i};\theta_{k}\right)}}+P_{k,n}\right\}\text{,}$
where $R_{k,n}\left(\theta_{k}\right)$ is the average negative log-likelihood the model in $\mathscr{H}_{k}$ defined by parameter $\theta_{k}\in\mathbb{T}_{k}$. This is often referred to as the problem of order selection in finite mixture modeling.
To fulfill B1 and B2, we can propose SWIC-type penalities of the form
$P_{k,n}$ $=\alpha q_{k}\sqrt{n^{-1}\log^{\left(\beta\right)}\left(n\right)}=\alpha k\left(1+q\right)\sqrt{n^{-1}\log^{\left(\beta\right)}\left(n\right)}\text{,}$ (7)
and in the case of B2, we can propose the typical BIC penalty
$P_{k,n}$ $=\frac{q_{k}}{2}\frac{\log n}{n}=\frac{k\left(1+q\right)}{2}\frac{\log n}{n}\text{.}$ (8)
Applying Theorem 1 is typically procedural, as A1 and A2 can also be checked directly. For example, supposing that we have a mixture of normal component densities:
$h_{k}\left(x;\theta_{k}\right)=\sum_{z=1}^{k}\pi_{k}\phi\left(x;\mu_{k},\sigma_{k}^{2}\right)\text{,}$
where
$\phi\left(x;\mu,\sigma^{2}\right)=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left\{-\frac{1}{2}\frac{\left(x-\mu\right)^{2}}{\sigma^{2}}\right\}$
is the normal density with mean $\mu\in\mathbb{R}$ and variance $\sigma^{2}>0$, $\theta_{k}=\left(\pi_{1},\dots,\pi_{k},\mu_{1},\sigma_{1}^{2},\dots,\mu_{k},\sigma_{k}^{2}\right)$, and $\mu_{k}\in\left[-\mathfrak{m},\mathfrak{m}\right]$ and $\sigma_{k}^{2}\in\left[1/\mathfrak{s},\mathfrak{s}\right]$ for constants $\mathfrak{m}>0$ and $\mathfrak{s}>1$. Then, we have the fact that $h\left(x;\cdot\right)$ is differentiable, for each $x\in\mathbb{R}$, with gradient:
$\frac{\partial\log h_{k}\left(x;\theta_{k}\right)}{\partial\theta_{k}}=\left(\dots,\frac{\phi\left(x;\mu_{k},\sigma_{k}^{2}\right)}{h\left(x;\theta_{k}\right)},\frac{x-\mu_{1}}{\sigma_{1}^{2}}\pi_{1}\frac{\phi\left(x;\mu_{1},\sigma_{1}^{2}\right)}{h\left(x;\theta_{k}\right)},\frac{\left(\mu_{1}-x\right)^{2}-\sigma_{1}^{2}}{2\sigma_{1}^{4}}\pi_{1}\frac{\phi\left(x;\mu_{1},\sigma_{1}^{2}\right)}{h\left(x;\theta_{k}\right)},\dots\right)$
and so $-\log h_{k}\left(x;\cdot\right)$ is Lipschitz with a constant of the form
$\mathfrak{C}\left(x\right)=k\times\max_{\mu\in\left[-\mathfrak{m},\mathfrak{m}\right],\sigma^{2}\in\left[1/\mathfrak{s},\mathfrak{s}\right]}\max\left\{1,\frac{\left|x-\mu\right|}{\sigma^{2}},\frac{\left(x-\mu\right)^{2}+\sigma^{2}}{2\sigma^{4}}\right\}\text{,}$
since $\phi\left(x;\mu_{z},\sigma_{z}^{2}\right)/h_{k}\left(x;\theta_{k}\right)$ and $\pi_{k}$ are bounded by 1, for each $z\in\left[k\right]$, via the mean value theorem. Then, we can verify that A3 holds by noting that normal random variables have finite moments of every order. A2 can be verified by the same observation.
For general sets $\mathbb{T}_{k}$, we cannot verify C1 since finite mixture models are only identifiable up to a permutation of the parameters. That is, if $-\log h\left(x;\cdot\right)$ is minimized at $\theta_{k}^{*}$, then it is also minimized at
$\theta_{k}^{\Pi}=\left(\pi_{\Pi\left(1\right)}^{*},\dots,\pi_{\Pi\left(k\right)}^{*},\mu_{\Pi\left(1\right)}^{*},\sigma_{\Pi\left(1\right)}^{*2},\dots,\mu_{\Pi\left(k\right)}^{*},\sigma_{\Pi\left(k\right)}^{*2}\right)\text{,}$
where $\Pi:\left[k\right]\rightarrow\left[k\right]$ is a permutation. However, typically, finite mixture models have isolated minima (6) and thus we can always consider $\mathbb{T}_{k}$ to be a compact set that contains only one such optimal parameter. C2 is also more onerous than A3, but is typically verified for common component densities. C3 and C4 can be difficult to verify, given that mixture models often have singular Fisher information (see, e.g. Watanabe, 2009, Sec. 7.8). Thus, it is difficult to make use of Theorem 2 for finite mixture models. However, this does not dismiss the use of BIC-like IC, as demonstrated by the specific analyses of Leroux (1992), Keribin (2000), and Gassiat (2018, Ch. 4), who establish conditions under which BIC-like IC can be proved consistent for mixture model order selection.
Remark 7.
We note that since $\left(\mathscr{H}_{k}\right)_{k\in\left[m\right]}$ is a nested sequence of hypotheses, our problem is only nontrivial if we assume that for sufficiently large $m$, there exists some $k^{*}<m$ such that $\inf_{\theta_{k^{*}}\in\mathbb{T}_{k^{*}}}r_{k^{*}}\left(\theta_{k^{*}}\right)=\inf_{\theta_{k}\in\mathbb{T}_{k}}r_{k}\left(\theta_{k}\right)$, for any $k>k^{*}$. This is true, for example, in the well-specified case, where the underlying data generating process is has density function in $\mathscr{H}_{k}$, for every $k\geq k^{*}$. That is, the data $\left(X_{i}\right)_{i\in\left[n\right]}$ are sampled IID from a normal mixture model with $k^{*}\in\left[m\right]$ components. If this is not the case, in the sense that $\inf_{\theta_{k}\in\mathbb{T}_{k}}r_{k}\left(\theta_{k}\right)$ is a strictly decreasing function of $k\in\mathbb{N}$, then the order selection problem above no longer makes sense and we then seek instead to choose a map $n\mapsto k\left(n\right)$, such that the loss expected loss $\text{E}\left[\inf_{\theta_{k\left(n\right)}\in\mathbb{T}_{k\left(n\right)}}R_{k\left(n\right),m}\left(\theta_{k\left(n\right)}\right)\right]$ converges to the limit $\lim_{k\rightarrow\infty}\inf_{\theta_{k}\in\mathbb{T}_{k}}r_{k}\left(\theta_{k}\right)$ at the fastest rate. For the minimum average negative log-likelihood problem, above, such results are considered in Li and Barron (1999) and Rakhlin et al. (2005). In particular, when the assumptions of Rakhlin et al. (2005) hold, the optimal rate of should take $k\left(n\right)=O\left(\sqrt{n}\right)$, which yields an optimal convergence rate of $O\left(1/\sqrt{n}\right)$. Examples of recent and more sophisticated results in this direction can be found in Ho et al. (2022).
3.2 Least absolute deviation and $\epsilon$-support vector regression
We suppose that we observe IID data $\left(X_{i}\right)_{i\in\left[n\right]}$, where each $X_{i}$ has the same distribution as $X=\left(Y,W\right)\in\mathbb{X}=\mathbb{Y}\times\mathbb{W}$, $\mathbb{Y}\subset\mathbb{R}$, and $\mathbb{W}\subset\mathbb{R}^{m}$. For each $k\in\left[m\right]$, we let
$\mathscr{H}_{k}=\left\{h_{k}\left(x;\theta_{k}\right)=y-\left\langle w^{\left(k\right)},\theta_{k}\right\rangle:\theta_{k}\in\mathbb{T}_{k}\right\}$ (9)
be the residual function obtained from the estimation of $y\in\mathbb{R}$ by $\left\langle w^{\left(k\right)},\theta_{k}\right\rangle$ (the inner product of $w^{\left(k\right)}$ and $\theta_{k}$), where $w^{\left(k\right)}=\left(w_{1},\dots,w_{k}\right)\in\mathbb{R}^{k}$ is the first $k$ elements of the vector $w\in\mathbb{R}^{m}$. Here, we assume that $\mathbb{T}_{k}\subset\mathbb{R}^{k}$ is compact, for each $k\in\left[m\right]$. We shall measure the goodness of the approximation $\left\langle w^{\left(k\right)},\theta_{k}\right\rangle$, for each $k$, using the $\epsilon$-insensitive $\text{L}_{1}$ loss
$\ell_{\epsilon}\left(y\right)=\left(\left|y\right|-\epsilon\right)_{+}\text{,}$
where $\epsilon\geq 0$ and $\left(\cdot\right)_{+}=\max\left\{0,\cdot\right\}$. When $\epsilon=0$, we simply have the $\text{L}_{1}$ loss: $\ell_{0}\left(y\right)=\left|y\right|$. For hypothesis class $\mathscr{H}_{k}$, we define
$\mathbb{T}_{k}^{*}=\underset{\theta_{k}\in\mathbb{T}}{\arg\min}\text{ }\underset{=r_{k}\left(\theta_{k}\right)}{\underbrace{\text{E}_{X}\left\{\ell_{\epsilon}\left(h_{k}\left(X;\theta_{k}\right)\right)\right\}}}$
to be the set of optimal linear $\epsilon$-support vector regression ($\epsilon$-SVR) parameters, with typical elements $\theta_{k}^{*}$, which we can estimate empirically by an $\epsilon$-SVR estimator (cf. Scholkopf and Smola, 2002, Sec. 9.1):
$\hat{\theta}_{k,n}\in\underset{\theta_{k}\in\mathbb{T}}{\arg\min}\text{ }\frac{1}{n}\sum_{i=1}^{n}\ell_{\epsilon}\left(h_{k}\left(X_{i};\theta_{k}\right)\right)\text{.}$ (10)
When $\epsilon=0$, (10) is commonly referred to as a least absolute deviation (LAD) estimator. Supposing that
$k^{*}=\min\underset{k\in\left[m\right]}{\arg\min}\left\{\min_{\theta_{k}\in\mathbb{T}_{k}}\text{E}_{X}\left\{\ell_{\epsilon}\left(h_{k}\left(X;\theta_{k}\right)\right)\right\}\right\}\text{,}$
which we wish to estimate. This is a version of the classical variable selection problem in regression analysis. The typical best-subset type variable selection problem (see, e.g., Heinze et al., 2018) can be obtained by instead considering $\mathscr{H}_{k}$ to include all models that depend on $k$ entries of $w$, rather than only the model that depends on the first $k$ entries of $w$ as per (9). One can then use the suggest IC to choose the optimal number of parameters $k^{*}$, and then within the models in $\mathscr{H}_{k^{*}}$, where all models are equally parsimonious, one can compare the respective sample minimum risks in order to decide which of the models with $k^{*}$ covariates is most optimal.
We can suggest IC for estimating $k^{*}$ of form
$\hat{K}_{n}=\min\underset{k\in\left[m\right]}{\arg\min}\text{ }\left\{\min_{\theta_{k}\in\mathbb{T}_{k}}\underset{=R_{k,n}\left(\theta_{k}\right)}{\underbrace{-\frac{1}{n}\sum_{i=1}^{n}\ell_{\epsilon}\left(h_{k}\left(X_{i};\theta_{k}\right)\right)}}+P_{k,n}\right\}\text{,}$ (11)
where
$P_{k,n}=\alpha k\sqrt{n^{-1}\log_{+}^{\left(\beta\right)}\left(n\right)}\text{and }P_{k,n}=\alpha kn^{-1}\log_{+}^{\left(\beta\right)}\left(n\right)\text{,}$ (12)
fulfill the requirements of B2 and B2, respectively.
Next, we observe that $\ell_{k}\left(x;\theta_{k}\right)=\ell_{\epsilon}\left(h_{k}\left(x;\theta_{k}\right)\right)$ is piecewise continuously differentiable in $\theta_{k}\in\mathbb{R}^{k}$, for each fixed $x\in\mathbb{R}^{k}$. Furthermore, this implies that it is therefore Lipschitz, with constant $\mathfrak{C}\left(x\right)=\mathfrak{C}\times\left\|w^{\left(k\right)}\right\|_{1}$, for some $\mathfrak{C}>0$, since the the gradient of $\ell_{k}\left(x;\theta_{k}\right)$, with respect to $\theta_{k}$, is either ±$\pm w^{\left(k\right)}$ or 00 (cf. Scholtes, 2012, Cor. 4.1.1). Thus, we only require that $\text{E}\left\|W\right\|_{2}^{2}$ is finite to verify A2. To further verify A1, we only additionally need that $\text{E}\left\{Y^{2}\right\}$ be finite.
For arbitrary measures on $X$ and sets $\mathbb{T}_{k}$, we cannot guarantee that each $r_{k}\left(\theta_{k}\right)$ has a unique minimizer. In fact, although trivial, one can construct examples whereupon $\mathbb{T}_{k}$ is uncountable and $\mathbb{T}_{k}^{*}=\mathbb{T}_{k}$. As an example, for $k=1$, suppose that $W^{\left(1\right)}=1$. Then, we can write
$r_{1}\left(\theta_{1}\right)=\text{E}_{X}\left(\left|Y-\theta_{1}\right|-\epsilon\right)_{+}\text{.}$
Now, suppose that $Y$ is supported on a bounded set $\mathbb{Y}\subset\mathbb{R}$, and that $\sup_{y\in\mathbb{Y},\theta_{1}\in\mathbb{T}_{1}}\left|y-\theta_{1}\right|<\epsilon$. Then, for every choice of $\theta_{1}\in\mathbb{T}_{1}$, $r_{1}\left(\theta_{1}\right)=0$. Thus, although somewhat pathological, it is possible to invalidate the uniqueness assumption in C1. Furthermore, we have also established that it is possible for the minimizers of $r_{k}\left(\theta_{k}\right)$ to be connected, thus unavailing the solution of localizing $\mathbb{T}_{k}$ around an appropriate isolated minimizer.
Of course, the construction above is an exclusive pathology to the $\epsilon>0$ case. Unfortunately, it is still not generally possible to guarantee uniqueness of the minimizer when $\epsilon=0$, without making additional assumptions. For example, in Pollard (1991), it is assumed that
$Y=\left\langle\theta_{k},W^{\left(k\right)}\right\rangle+E\text{,}$
where $E$ is a random variable with median zero, and a continuous and positive density function in a neighborhood of zero, say, to guarantee that C1 holds for $k\in\left[m\right]$. A similar approach is taken in Knight (1998), who assume differentiability at zero of the distribution function of the noise variable $E$. The approaches of Pollard (1991) and Knight (1998) make strong assumptions regarding the data generating process, but both provide asymptotic normality of the risk minimizer $\theta_{k}^{*}$, if it is unique. Once obtained, a delta method or asymptotic expansion approach (such as in Sin and White (1996)) can be used to obtain the boundedness in probability of the $\sqrt{n}$ or $n$ scaling of the minimum risk. This differs to our approach, which seeks to obtain the asymptotic law for the $\sqrt{n}$ scaling of the minimum risk, directly.
We notice that if $X$ has measure that is absolutely continuous with respect to the Lebesgue measure on $\mathbb{X}$, then $\ell_{k}\left(X;\theta_{k}\right)=\left(\left|Y-\left\langle\theta_{k},W^{\left(k\right)}\right\rangle\right|-\epsilon\right)_{+}$ is twice differentiable with respect to $\theta_{k}\in\mathbb{T}_{k}$, for almost all $X\in\mathbb{X}$, for each $k\in\left[m\right]$ but that the second derivative will be zero for every $\theta_{k}$. This invalidates the necessary and sufficient condition for C3 under C4, as per Appendix A, and thus we cannot verify the consistency of BIC-like criteria using Theorem 2. Nevertheless, at least in the $\epsilon=0$ case, BIC-like criteria have been proved consistent for variable selection under various assumptions (see, e.g., Bai, 1998 and Lee et al., 2014).
Finally, we note that our construction above is a simplified version of the variable selection problem. The typical best-subset type variable selection problem (see, e.g., Heinze et al., 2018) can be obtained by instead considering $\mathscr{H}_{k}$ to include all models that depend on $k$ entries of $w$, rather than only the model that depends on the first $k$ entries of $w$ as per (9). One can use the suggest IC to choose the optimal number of parameters $k^{*}$, and then within the models in $\mathscr{H}_{k^{*}}$, where all models are equally parsimonious, one can compare the respective sample minimum risks in order to decide which of the models with $k^{*}$ covariates is most optimal.
3.3 Principal component analysis
We next consider $\left(X_{i}\right)_{i\in\left[n\right]}$ IID with each $X_{i}$ from having the same distribution as the random variable $X\in\mathbb{X}\subset\mathbb{R}^{m}$. For each $k\in\left[m\right]$, we let
$\mathscr{H}_{k}=\left\{h_{k}\left(x;\theta_{k}\right)=x-\theta_{k}\theta_{k}^{\top}x:\theta_{k}\in\mathbb{T}_{k}\right\}\text{,}$ (13)
where $\mathbb{T}_{k}\subset\mathbb{R}^{m\times k}$ and $\left(\cdot\right)^{\top}$ is the transposition operation. Like in Section 3.2, we can consider $h_{k}\left(x;\theta_{k}\right)$ as the residual function under a $k\leq m$ dimensional linear subspace encoding and decoding operations defined via $\theta_{k}^{\top}$ and $\theta_{k}$, respectively. If we choose the quadratic loss $\ell\left(\cdot\right)=\left\|\cdot\right\|_{2}^{2}$, then
$\underset{\theta_{k}\in\mathbb{T}_{k}}{\arg\min}\underset{=R_{k,n}\left(\theta_{k}\right)}{\underbrace{\text{ }\frac{1}{n}\sum_{i=1}^{n}\left\|X_{i}-\theta_{k}\theta_{k}^{\top}X_{i}\right\|_{2}^{2}}}$ (14)
is the classical problem of computing the first $k$ principal components of the data $\left(X_{i}\right)_{i\in\left[n\right]}$, and obtaining solutions to (14) is conventionally referred to as principal component analysis (PCA). We can conveniently choose any compact parameter space $\mathbb{T}_{k}$ containing $\mathbb{O}_{k}=\left\{\theta_{k}\in\mathbb{R}^{m\times k}:\theta_{k}^{\top}\theta_{k}=\mathbf{I}_{k}\right\}$, where $\mathbf{I}_{k}$ is the identity of $\mathbb{R}^{k\times k}$, since
$\underset{\theta_{k}\in\mathbb{R}^{m\times k}}{\arg\min}\text{ }\frac{1}{n}\sum_{i=1}^{n}\left\|X_{i}-\theta_{k}\theta_{k}^{\top}X_{i}\right\|_{2}^{2}=\underset{\theta_{k}\in\mathbb{O}_{k}}{\arg\min}\text{ }\frac{1}{n}\sum_{i=1}^{n}\left\|X_{i}-\theta_{k}\theta_{k}^{\top}X_{i}\right\|_{2}^{2}$
via Shalev-Shwartz and Ben-David (2014, Lem. 23.1).
We can write
$\ell_{k}\left(x;\theta_{k}\right)=\left\|x-\theta_{k}\theta_{k}^{\top}x\right\|_{2}^{2}=\left\|x\right\|_{2}^{2}-\text{trace}\left\{\theta_{k}^{\top}xx^{\top}\theta_{k}\right\}\text{,}$ (15)
which yields the expression
$r_{k}\left(\theta_{k}\right)=\text{trace}\left\{\text{E}_{X}\left(XX^{\top}\right)\right\}-\text{trace}\left\{\theta_{k}^{\top}\text{E}_{X}\left(XX^{\top}\right)\theta_{k}\right\}\text{,}$ (16)
via the linearity of the trace. Clearly, when $k=m$, we can choose $\theta_{k}^{*}=\text{{I}}_{m}$ to obtain the minimum value $r_{k}\left(\theta_{k}^{*}\right)=0$. However, suppose that we wish to search for the smallest $k^{*}<m$ such that $r_{k^{*}}\left(\theta_{k^{*}}^{*}\right)=0$. Then, we may identify such an optimal hypothesis index with the optimization problem
$k^{*}=\min\underset{k\in\left[m\right]}{\arg\min}\left\{\min_{\theta_{k}\in\mathbb{T}_{k}}r_{k}\left(\theta_{k}\right)\right\}\text{,}$
which we can then estimate via the IC
$\hat{K}_{n}=\min\underset{k\in\left[m\right]}{\arg\min}\text{ }\left\{\min_{\theta_{k}\in\mathbb{T}_{k}}\text{ }\frac{1}{n}\sum_{i=1}^{n}\left\|X_{i}-\theta_{k}\theta_{k}^{\top}X_{i}\right\|_{2}^{2}+P_{k,n}\right\}\text{,}$ (17)
where we can take $P_{k,n}$ the same as (12) to satisfy B2 or B2.
Since (15) is quadratic, we can then compute the derivative with respect to $\theta_{k}$ to be $\partial\ell_{k}\left(x;\theta_{k}\right)/\partial\theta_{k}=2xx^{\top}\theta_{k}$. We can thus verify A2 by assuming that the fourth moment $\text{E}_{X}\left\|X\right\|_{4}^{4}$ is finite, which then also verifies A1.
Noting the form of (16), we then conclude, by Shalev-Shwartz and Ben-David (2014, Thm. 23.2), that $r_{k}$ is minimized at any
$\theta_{k}^{*}=\left[\begin{array}[]{ccc}\upsilon_{\Pi\left(1\right)}&\cdots&\upsilon_{\Pi\left(k\right)}\end{array}\right]\text{,}$ (18)
where $\upsilon_{1},\dots,\upsilon_{k}\in\mathbb{R}^{m}$ are orthonormal eigenvectors corresponding to the first $k$ largest eigenvalues of $\text{E}_{X}\left(XX^{\top}\right)$ and $\Pi$, again, is a permutation of $\left[k\right]$. Thus, C1 does not hold unless we localize $\mathbb{T}_{k}$ on a neighborhood of one of the isolated matrices. C2 is trivial, and C3 and C4 can then be verified when $\text{E}_{X}\left(XX^{\top}\right)$ is positive definite, using the results from Appendix A. We note that conditions under which the BIC is consistent for determining the optimal subspace dimension of data has been previously considered in the works of Zhao et al. (1986), Fujikoshi and Sakurai (2016), and Bai et al. (2018).
4 Numerical experiments
The discussions thus far have been concerned with the theoretical performance of PanIC and BIC-like estimators when sample sizes approach infinity. However, practical behaviors of such estimators vary when sample sizes are small. As such, we find it useful to compare criteria in small sample settings to examine their relative performances.
For a strictly increasing sequence of constants $\left(c_{k}\right)_{k\in\left[m\right]}$, we will say that the (general) BIC is an estimator of form (2), where we choose $P_{n,k}=c_{k}n^{-1}\log n/2$, for each $k\in\left[m\right]$. Since the SWIC defined in Remark 3 require a choices for the parameters $\alpha$ and $\beta$, we find it practical to couple the choice of $\alpha$ with the BIC, by choosing constants $\alpha=\alpha\left(\beta,\nu\right)>0$, which are defined as the value of $\alpha$ that makes the order $\beta$ SWIC and the BIC equal, when the sample size is $\nu\in\mathbb{N}$:
$\alpha\left(\beta,\nu\right)=\frac{\log\nu}{2\sqrt{\nu\log_{+}^{\left(\beta\right)}\left(\nu\right)}}\text{.}$
To facilitate the comparison of the orders $\beta\in\left\{1,2\right\}$ SWIC with the BIC, we will consider the respective choices of $\alpha\left(\beta,\nu\right)$, with $\nu\in\left\{10^{3},10^{4}\right\}$. We will also provide comparisons with the AIC ($P_{n,k}=c_{k}/n$), for completeness.
We conduct all experiments in the R programming language and make our code public at: https://github.com/hiendn/PanIC.
4.1 Normal finite mixture models
In this experiment, we simulate data $\left(X_{i}\right)_{i\in\left[n\right]}$ from bivariate normal mixture models with densities of forms
$\sum_{z=1}^{k^{*}}\pi_{z}\phi\left(x;\mu_{z},\Sigma_{z}\right)\text{,}$
where $\phi\left(\cdot;\mu,\Sigma\right)$ is the normal density with mean vector $\mu\in\mathbb{R}^{2}$ and covariance matrix $\Sigma\in\mathbb{R}^{2\times 2}$. Here, we consider four different scenarios:
  • S1.1
    $k^{*}=4$ and
    $\left(\mu_{z}\right)_{z\in\left[4\right]}=2\times\left\{\left[\begin{array}[]{c}1\\ 1\end{array}\right],\left[\begin{array}[]{c}1\\ 2\end{array}\right],\left[\begin{array}[]{c}2\\ 1\end{array}\right],\left[\begin{array}[]{c}2\\ 2\end{array}\right]\right\}\text{.}$
  • S1.2
    $k^{*}=4$ and
    $\left(\mu_{z}\right)_{z\in\left[4\right]}=3\times\left\{\left[\begin{array}[]{c}1\\ 1\end{array}\right],\left[\begin{array}[]{c}1\\ 2\end{array}\right],\left[\begin{array}[]{c}2\\ 1\end{array}\right],\left[\begin{array}[]{c}2\\ 2\end{array}\right]\right\}\text{.}$
  • S1.3
    $k^{*}=6$ and
    $\left(\mu_{z}\right)_{z\in\left[6\right]}=2\times\left\{\left[\begin{array}[]{c}1\\ 1\end{array}\right],\left[\begin{array}[]{c}1\\ 2\end{array}\right],\left[\begin{array}[]{c}1\\ 3\end{array}\right],\left[\begin{array}[]{c}2\\ 1\end{array}\right],\left[\begin{array}[]{c}2\\ 2\end{array}\right],\left[\begin{array}[]{c}2\\ 3\end{array}\right]\right\}\text{.}$
  • S1.4
    $k^{*}=6$ and
    $\left(\mu_{z}\right)_{z\in\left[6\right]}=3\times\left\{\left[\begin{array}[]{c}1\\ 1\end{array}\right],\left[\begin{array}[]{c}1\\ 2\end{array}\right],\left[\begin{array}[]{c}1\\ 3\end{array}\right],\left[\begin{array}[]{c}2\\ 1\end{array}\right],\left[\begin{array}[]{c}2\\ 2\end{array}\right],\left[\begin{array}[]{c}2\\ 3\end{array}\right]\right\}\text{.}$
In all scenarios S1.1–S1.4, we set $\pi_{z}=1/k^{*}$ and $\Sigma_{z}=\mathbf{I}_{2}$, for each $z\in\left[k^{*}\right]$.
We take a sequence of hypothesis of form (5) with $m=8$ and assess the performances of the AIC, BIC and SWIC of order $\beta\in\left\{1,2\right\}$, using penalities of forms $k\left(q+1\right)/n$, (8), and (7), respectively, with $q=5$. To evaluate performances, we consider sample sizes $n\in\left\{10^{2},10^{3},10^{4},10^{5}\right\}$ and apply the IC on 100 simulation runs, and compute the average value of $\hat{K}_{n}$ obtained via each IC (Avg), as well as the proportion of occasions when $\hat{K}_{n}=k^{*}$ (Prop). The experimental results are presented in Tables 1. Per Remark 7, we note that the order selection procedure is nontrivial since we are operating in the well-specified parametric setting where the data generating process of $\left(X_{i}\right)_{i\in\left[n\right]}$ corresponds to one of the hypotheses in $\left(\mathscr{H}_{k}\right)_{k\in\left[m\right]}$.
Table 1: Results from 100 simulation runs for each scenario of the normal finite mixture models experiment. Underlined values highlight either that the corresponding IC has either the closest Avg to $k^{*}$ or the greatest Prop value for the indicated scenario.
SWIC $\left(\beta,\nu\right)$
Scenario $n$ AIC BIC $\left(1,10^{3}\right)$ $\left(1,10^{4}\right)$ $\left(2,10^{3}\right)$ $\left(2,10^{4}\right)$
Avg Prop Avg Prop Avg Prop Avg Prop Avg Prop Avg Prop
S1.1 $10^{2}$ 1.54 0.02 1.00 0.00 1.77 0.03 6.48 0.08 1.55 0.02 6.12 0.08
$\left(k^{*}=4\right)$ $10^{3}$ 2.28 0.08 1.03 0.00 1.03 0.00 1.92 0.02 1.03 0.00 1.83 0.01
$10^{4}$ 4.02 0.88 3.32 0.32 2.19 0.00 3.32 0.32 2.36 0.00 3.32 3.32
$10^{5}$ 5.02 0.28 4.31 0.71 3.39 0.39 4.07 0.89 3.63 0.63 4.09 0.87
S1.2 $10^{2}$ 2.66 0.09 1.06 0.00 2.97 0.14 6.86 0.10 2.72 0.10 6.52 0.11
$\left(k^{*}=4\right)$ $10^{3}$ 4.05 0.97 3.53 0.54 3.53 0.54 4.00 1.00 3.53 0.54 4.00 1.00
$10^{4}$ 4.00 1.00 4.00 1.00 4.00 1.00 4.00 1.00 4.00 1.00 4.00 1.00
$10^{5}$ 4.00 1.00 4.00 1.00 4.00 1.00 4.00 1.00 4.00 1.00 4.00 1.00
S1.3 $10^{2}$ 1.45 0.01 1.00 0.00 1.69 0.04 6.84 0.14 1.47 0.01 6.00 0.10
$\left(k^{*}=6\right)$ $10^{3}$ 2.73 0.00 1.67 0.00 1.67 0.00 2.57 0.00 1.67 0.00 2.53 0.00
$10^{4}$ 4.91 0.20 3.84 0.00 2.37 0.00 3.84 0.00 2.44 0.00 3.84 0.00
$10^{5}$ 7.04 0.20 5.89 0.45 3.96 0.00 4.64 0.11 4.01 0.00 4.67 0.12
S1.4 $10^{2}$ 2.84 0.04 1.07 0.00 3.30 0.06 6.59 0.21 3.02 0.05 6.26 0.23
$\left(k^{*}=6\right)$ $10^{3}$ 5.64 0.53 3.61 0.01 3.61 0.01 5.30 0.43 3.61 0.01 5.20 0.40
$10^{4}$ 6.11 0.90 6.09 0.91 5.75 0.84 6.09 0.91 5.87 0.89 6.09 0.91
$10^{5}$ 6.18 0.85 6.14 0.87 6.05 0.95 6.05 0.95 6.05 0.95 6.05 0.95
We firstly notice that no IC is dominant across the range of scenarios and sample sizes. However, we can make a number of generalization. Firstly, the choices of $\alpha$ constants for the SWIC leads to criteria that are highly anti-conservative when $n=100$, in the sense that the estimates $\hat{K}_{n}$ tend to be larger than the BIC and even the AIC. This is typically a bad thing, since one generally would like to choose the most parsimonious model, but we can observe that this propensity to choose larger hypothesis classes can lead to greater proportions of correct identifications of $k^{*}$.
This anti-conservativeness decreases dramatically, when comparing the results for $n=1000$ to $n=100$, where we observe that the SWIC become more conservative than the AIC, where the constants based on $\nu=10^{4}$ tend to resemble the performance of the AIC in all scenarios.
The AIC appears to perform well across all scenarios when $10^{4}$, although, as expected by the fact that the AIC satisfies B1 but not B2 (or B2), when $n$ is further increased to $10^{5}$, we observe that the AIC becomes anti-conservative in all cases except for S1.2, where all IC perform perfectly. Most importantly, we observe that, as expected from Theorem 1, regardless of the choice of $\alpha$ and $\beta$, the Avg values for the SWIC appear to converge towards to $k^{*}$, as $n$ increases in all settings, albeit rather slowly in the case of S1.3. We also observe the same convergence regarding the BIC, which supports the theory of Leroux (1992), Keribin (2000), and Gassiat (2018, Ch. 4).
4.2 Least absolute deviation and $\epsilon$-support vector regression
Let $X=\left(Y,W\right)\in\mathbb{R}\times\mathbb{R}^{m}$, for $m=10$. We simulate $\left(X_{i}\right)_{i\in\left[n\right]}$, IID, from the same distribution as $X$, defined via the linear relationship
$Y=\left\langle\theta^{*},W\right\rangle+E\text{,}$
where $W\sim\text{Uniform}\left(\left[0,1\right]^{m}\right)$ and $E\sim\text{N}\left(0,1\right)$. Here, we consider four scenarios S2.1–S2.4, whereupon in S2.1 and S2.2,
$\theta^{*}=\left(1,0.75,0.5,0.25,0,0,0,0,0,0\right)$
and in S2.3 and S2.4,
$\theta^{*}=\left(1,0.85,0.7,0.55,0.4,0.25,0,0,0,0\right)\text{.}$
Using the sequence of hypotheses of for(9) with $m=10$, we assess the performances of the AIC, BIC and SWIC of order $\beta\in\left\{1,2\right\}$ of form (11) with penalties $k/n$, $kn^{-1}\log n/2$, and $\alpha k\sqrt{n^{-1}\log_{+}^{\left(\beta\right)}\left(n\right)}$, respectively. Here, $\epsilon=0$ in S2.1 and S2.3, and $\epsilon=1$ in S2.2 and S2.4. We further note that $k^{*}=4$ in S2.1 and S2.2, and $k^{*}=6$ in S2.3 and S2.4. To evaluate the performances of the IC, we consider sample sizes $n\in\left\{10^{2},10^{3},10^{4},10^{5}\right\}$ and use 100 simulation runs of each scenario. Avg and Prop values are recorded in Table 2.
Table 2: Results from 100 simulation runs for each scenario of the LAD and SVR experiment. Underlined values highlight either that the corresponding IC has either the closest Avg to $k^{*}$ or the greatest Prop value for the indicated scenario.
SWIC $\left(\beta,\nu\right)$
Scenario $n$ AIC BIC $\left(1,10^{3}\right)$ $\left(1,10^{4}\right)$ $\left(2,10^{3}\right)$ $\left(2,10^{4}\right)$
Avg Prop Avg Prop Avg Prop Avg Prop Avg Prop Avg Prop
S2.1 $10^{2}$ 3.62 0.16 2.30 0.01 4.18 0.14 7.97 0.07 3.66 0.16 7.63 0.08
$\left(k^{*}=4\right)$ $10^{3}$ 4.47 0.52 3.39 0.37 3.39 0.37 4.23 0.51 3.39 0.37 4.12 0.54
$\left(\epsilon=0\right)$ $10^{4}$ 5.13 0.60 4.02 0.98 3.98 0.98 4.02 0.98 3.99 0.99 4.02 0.98
$10^{5}$ 4.75 0.69 4.00 1.00 4.00 1.00 4.00 1.00 4.00 1.00 4.00 1.00
S2.2 $10^{2}$ 2.69 0.05 1.95 0.00 2.76 0.06 6.06 0.13 2.70 0.05 5.34 0.16
$\left(k^{*}=4\right)$ $10^{3}$ 3.71 0.61 3.13 0.17 3.13 0.17 3.59 0.57 3.13 0.17 3.55 0.55
$\left(\epsilon=1\right)$ $10^{4}$ 4.17 0.89 3.99 0.99 3.79 0.79 3.99 0.99 3.87 0.87 3.99 0.99
$10^{5}$ 4.22 0.92 4.00 1.00 4.00 1.00 4.00 1.00 4.00 1.00 4.00 1.00
S2.3 $10^{2}$ 4.94 0.14 3.45 0.01 5.24 0.17 7.79 0.09 5.08 0.16 7.56 0.08
$\left(k^{*}=6\right)$ $10^{3}$ 6.76 0.55 5.37 0.38 5.37 0.38 6.39 0.58 5.37 0.38 6.31 0.60
$\left(\epsilon=0\right)$ $10^{4}$ 6.74 0.64 6.02 0.98 5.96 0.96 6.02 0.98 5.97 0.97 6.02 0.98
$10^{5}$ 6.79 0.65 6.02 0.98 6.00 1.00 6.00 1.00 6.00 1.00 6.00 1.00
S2.4 $10^{2}$ 3.91 0.09 2.92 0.00 4.20 0.08 6.78 0.13 3.95 0.09 6.29 0.15
$\left(k^{*}=6\right)$ $10^{3}$ 5.64 0.50 4.89 0.12 4.89 0.12 5.42 0.41 4.89 0.12 5.36 0.38
$\left(\epsilon=1\right)$ $10^{4}$ 6.14 0.91 6.00 1.00 5.64 0.64 6.00 1.00 5.80 0.80 6.00 1.00
$10^{5}$ 6.10 0.91 6.00 1.00 6.00 1.00 6.00 1.00 6.00 1.00 6.00 1.00
As in Section 4.1, we observe that when $n=100$, the SWIC with the considered choices for $\alpha$ can be highly anti-conservative, especially when $\nu=10000$. This can lead to over-estimation of $k^{*}$, although this can sometimes lead to a higher value of Prop. For $n>100$, we observe that the SWIC and BIC all tend to perform similarly, and in all cases the Avg and Prop appear to converge, to $k^{*}$ and $1$, respectively, as $n$ increases. In the case of the SWIC, this is predictable from Theorem 1, and the consistency of the BIC is in the LAD scenarios (S2.1 and S2.3) is also supported by the works of Bai (1998) and Lee et al. (2014). However, the apparent consistency of the BIC in $\epsilon>0$ SVR scenarios is of interest and can likely be established via the same analysis as that of Lee et al. (2014). Proving this result is outside of the scope the current work. We finally note that the behaviors of the SWIC and BIC contrast with that of the AIC, which appears to remain anti-conservative, even when $n$ is large, in all scenarios.
4.3 Principal component analysis
We now simulate $\left(X_{i}\right)_{i\in\left[n\right]}$, IID, with the same distribution as $X\in\mathbb{R}^{m}$ with $m=10$, where
$X=\mathbf{A}Y$
for some fixed matrix $\mathbf{A}\in\mathbb{R}^{m\times k^{*}}$ and $k^{*}$-dimensional random variable $Y$ that differs in distribution according to the simulation scenario. Here, we take $Y\sim\text{N}\left(0,\mathbf{R}\right)$ in Scenarios S3.1 and S3.3, and $Y\sim t_{5}\left(0,\mathbf{R}\right)$ in S3.2 and S3.4, where $\mathbf{R}$ has diagonal elements equal to 1 and off-diagonal elements equal to $3/4$, and where $t_{\text{df}}$ denotes the multivariate Student-$t$ law with $\text{df}\in\mathbb{R}_{>0}$ degrees of freedom (cf. Fang et al. 1990, Sec. 3.3.6). We consider two different choices for $\mathbf{A}$:
$\left[\begin{array}[]{cccc}1&0&0&0\\ 0&0.25&0&0\\ 0&0&0.1&0\\ 0&0&0&0.1\\ 1&0.25&0&0\\ 0&0.25&0.1&0\\ 0&0&0.1&0.1\\ 1&0.25&0.1&0\\ 0&0.25&0.1&0.1\\ 1&0.25&0.1&0.1\end{array}\right]\left(\text{S3.1, S3.2: }k^{*}=4\right)\text{,}\left[\begin{array}[]{cccccc}1&0&0&0&0&0\\ 0&0.5&0&0&0&0\\ 0&0&0.25&0&0&0\\ 0&0&0&0.25&0&0\\ 0&0&0&0&0.1&0\\ 0&0&0&0&0&0.1\\ 1&0.5&0&0&0&0\\ 0&0.5&0.25&0&0&0\\ 0&0&0.25&0.25&0&0\\ 0&0&0&0.25&0.1&0\end{array}\right]\text{$\left(\text{S3.3, S3.4: }k^{*}=6\right)$.}$
Using the sequence of hypotheses of form (9) with $m=10$, we assess the performances of the AIC, BIC and SWIC of order $\beta\in\left\{1,2\right\}$ of form (11) with penalities $km/n$, $kmn^{-1}\log n/2$, and $\alpha km\sqrt{n^{-1}\log_{+}^{\left(\beta\right)}\left(n\right)}$, respectively. We again assess the IC via 100 simulation runs of each scenario with sample sizes $n\in\left\{10^{2},10^{3},10^{4},10^{5}\right\}$, and provide Avg and Prop values in Table (3).
Table 3: Results from 100 simulation runs for each scenario of the PCA experiment. Underlined values highlight either that the corresponding IC has either the closest Avg to $k^{*}$ or the greatest Prop value for the indicated scenario.
SWIC $\left(\beta,\nu\right)$
Scenario $n$ AIC BIC $\left(1,10^{3}\right)$ $\left(1,10^{4}\right)$ $\left(2,10^{3}\right)$ $\left(2,10^{4}\right)$
Avg Prop Avg Prop Avg Prop Avg Prop Avg Prop Avg Prop
S3.1 $10^{2}$ 2.00 0.00 1.00 0.00 2.00 0.00 2.99 0.00 2.00 0.00 2.90 0.00
$\left(k^{*}=4\right)$ $10^{3}$ 3.00 0.00 3.00 0.00 3.00 0.00 3.00 0.00 3.00 0.00 3.00 0.00
$\left(Y\sim\text{N}\right)$ $10^{4}$ 4.00 1.00 4.00 1.00 3.00 0.00 4.00 1.00 3.00 0.00 4.00 1.00
$10^{5}$ 4.00 1.00 4.00 1.00 4.00 1.00 4.00 1.00 4.00 1.00 4.00 1.00
S3.2 $10^{2}$ 2.11 0.00 1.72 0.00 2.24 0.00 3.00 0.00 2.14 0.00 3.00 0.00
$\left(k^{*}=4\right)$ $10^{3}$ 3.24 0.24 3.00 0.00 3.00 0.00 3.00 0.00 3.00 0.00 3.00 0.00
$\left(Y\sim t_{5}\right)$ $10^{4}$ 4.00 1.00 4.00 1.00 3.00 0.00 4.00 1.00 3.00 0.00 4.00 1.00
$10^{5}$ 4.00 1.00 4.00 1.00 4.00 1.00 4.00 1.00 4.00 1.00 4.00 1.00
S3.3 $10^{2}$ 2.08 0.00 1.26 0.00 2.20 0.00 3.10 0.00 2.12 0.00 3.00 0.00
$\left(k^{*}=6\right)$ $10^{3}$ 4.00 0.00 3.00 0.00 3.00 0.00 4.00 0.00 3.00 0.00 4.00 0.00
$\left(Y\sim\text{N}\right)$ $10^{4}$ 6.00 1.00 4.97 0.00 4.00 0.00 4.97 0.00 4.00 0.00 4.97 0.00
$10^{5}$ 6.00 1.00 6.00 1.00 5.00 0.00 6.00 1.00 5.00 0.00 6.00 1.00
S3.4 $10^{2}$ 2.89 0.00 1.98 0.00 2.99 0.00 3.89 0.00 2.93 0.00 3.65 0.00
$\left(k^{*}=6\right)$ $10^{3}$ 4.00 0.00 4.00 0.00 4.00 0.00 4.00 0.00 4.00 0.00 4.00 0.00
$\left(Y\sim t_{5}\right)$ $10^{4}$ 6.00 1.00 5.83 0.83 4.00 0.00 5.83 0.83 4.00 0.00 5.83 0.83
$10^{5}$ 6.00 1.00 6.00 1.00 6.00 1.00 6.00 1.00 6.00 1.00 6.00 1.00
We again observe that the SWIC tend to be anti-conservative when $n=100$, although we do not observe over-estimation of $k^{*}$ in any of our scenarios, for any of the criteria. This is especially interesting when inspecting the AIC results, are typically anti-conservative and not consistent. When $n>100$ the SWIC and BIC appear to behave similarly, with a tendency for Avg and Prop to approach $k^{*}$ and 1, respectively, in accordance to Theorems 1 and 2. We particularly note that $\nu=1000$ appears to be too conservative when choosing $\alpha$, as in S2.3, both of the SWIC with $\nu=1000$ do not correctly identify $k^{*}$, even when $n=10^{5}$. This conservativeness is further demonstrated when observing the results for $n=10^{4}$ in S3.1, S3.2 and S3.4, where the BIC appears to estimate $k^{*}$ correctly with high probability, but where the SWIC with $\nu=1000$ yield under-estimates of $k^{*}$.
5 Discussion
Model selection is a ubiquitous task that arises when using many methods for statistical inference and machine learning. The method of IC has long been an essential tool for model selection in the likelihood setting and related paradigms, such as quasi-likelihood and composite likelihood-based methods. Using asymptotic theory from the stochastic programming literature, we have demonstrated that simple and practical IC-based methods are available for conducting variable selection in general settings, such as when there are no clear definitions of likelihood functions, and when estimators may not be unique. Following the primary work of Sin and White (1996), we propose the PanIC framework for model selection, that requires minimal assumptions for guaranteeing consistent identification of parsimonious models. We also provide relaxations of sufficient conditions for proving consistency of BIC-like criteria, when compared to Sin and White (1996) and Baudry (2015). Our empirical studies demonstrate that PanIC estimators can be usefully applied in typical model selection settings where the consistency of the BIC-like criteria are hard to verify, or when there are no IC-based methods currently available.
Although general and useful, we have identified some directions to improve upon the present work. Firstly, while restrictive, sufficient conditions from Sin and White (1996) permit the possibility of non-IID data, when verifying the consistency of PanIC and BIC-like criteria, via generic non-IID strong laws of large numbers. Although non-trivial, inspections of the proofs of Lemmas 1 and 3 from Shapiro et al. (2021) reveal a reliance on empirical processes methods for proving Donsker-type central limit theorems that can be replaced by non-IID variants, such as those of Dedecker and Louhichi (2002) and Rio (2017, Ch. 8). Progress in this direction has been made by Wang et al. (2022), although their results are insufficient for establishing non-IID versions of Lemmas 1 and 3.
Secondly, one may contend that our choices of $\alpha$ for the SWIC in Section 4 are arbitrary and unlikely to be optimal, although we note that in the context of Theorem (1), the choice of $\alpha$ is immaterial when considering the asymptotic properties of the SWIC and other PanIC. However, in more critical applied work, it is possible to select $\alpha$ in a principled and discerning manner. One such approach would be to conduct pilot simulations of potential models from which data may arise and then apply the so-called slope heuristics of Birge and Massart (2007) and Arlot (2020) to estimate the value of $\alpha$. Such an approach has been usefully applied, for example, in Nguyen et al. (2018) and Nguyen et al. (2022b), via the software of Baudry et al. (2012).
Thirdly, we note that all of our results are asymptotic and thus provide no guarantees for any fixed sample size $n\in\mathbb{N}$. Finite sample oracle inequalities for estimators of form (2) have been comprehensively studied in the works of Massart (2007) and Giraud (2022), for example. Although desirable, the establishment of oracle inequalities may be excessive and require more bespoke and precise analyses that go beyond the simple verification of the sufficient conditions required by PanIC. We note that when $\left(\mathscr{H}_{k}\right)_{k\in\left[m\right]}$ is a sequence of nested hypotheses and when $\ell$ corresponds to a likelihood, composite likelihood, or conditional likelihood maximization problem, then the approach of Nguyen et al. (2022a) provides a straightforward method for obtaining finite sample confidence statements regarding the event $\left\{\omega:\tilde{K}_{n}\geq k^{*}\right\}$, where $\tilde{K}_{n}$ is an estimator of $k^{*}$ that is obtained via the universal inference hypothesis testing framework of Wasserman et al. (2020). Such confidence statements can be used in combination with PanIC to provide complementary measures of uncertainty regarding the value of $k^{*}$.
Lastly, although we have only concerned ourselves with sufficient conditions that guarantee (1), some situations may require the guarantee of strong consistency:
$\text{Pr}\left(\lim_{n\rightarrow\infty}\hat{K}_{n}=k^{*}\right)=1\text{.}$ (19)
Demanding regularity conditions are provided in Sin and White (1996) for proving (19) in the general setting, while setting-specific requirements are provided in the works of Zhao et al. (1986), Keribin (2000), and Gassiat (2018, Ch. 4), for example. This prompts the question as to whether there is a version of Theorem 1, with comparable assumptions, that guarantees (19). We are happy to provide an affirmative answer, thanks to the recent work of Banholzer et al. (2022). The following result can be inferred from Banholzer et al. (2022, Thm. 4).
Lemma 2.
Assume that $\left(X_{i}\right)_{i\in\left[n\right]}$ is an IID sequence, and that A1 and A2 hold, for each $k\in\left[m\right]$. Then, for each $k$,
$\sqrt{\frac{n}{\log_{+}^{\left(2\right)}\left(n\right)}}\left(\min_{\theta_{k}\in\mathbb{T}_{k}}R_{k,n}\left(\theta_{k}\right)-\min_{\theta_{k}\in\mathbb{T}_{k}}r_{k}\left(\theta_{k}\right)\right)=O_{\text{a.s.}}\left(1\right)\text{.}$ (20)
Make the following assumptions, for each $k\in\left[m\right]$:
B1
$P_{k,n}>0$, for each $k\in\left[m\right]$ and $n\in\mathbb{N}$, and $P_{k,n}=o_{\text{a.s.}}\left(1\right)$, as $n\rightarrow\infty$.
B2∗∗
If $k<l$, then $\sqrt{n/\log_{+}^{\left(2\right)}\left(n\right)}\left\{P_{l,n}-P_{k,n}\right\}\rightarrow\infty$, almost surely, as $n\rightarrow\infty$.
Theorem 3.
Assume that $\left(X_{i}\right)_{i\in\left[n\right]}$ is an IID sequence, and that A1, A2, B1 and B2∗∗ hold, for each $k\in\left[m\right]$. Then, the estimator (2) satisfies the strong consistency property (19).
Theorem 3 can be considered as a strong consistency version of Theorem 1. We may refer to the class of IC that satisfies B1 and B2∗∗ as strong PanIC. Note that SWIC of order 1 are strong PanIC but SWIC with $\beta>1$ and the BIC are not. This provides some incentive towards the use of the order 1 PanIC over those of higher order, although we observe from Section 4 that the difference may not be so pronounced in finite samples.
We note that the BIC and Hannan–Quinn IC can be proved to be strongly consistent in many settings, but do not satisfy the assumptions of Theorem 3. Unfortunately, we cannot yet prove a strong consistency equivalent to Theorem 2 that would apply to the BIC and Hannan–Quinn IC. We believe that such a result is available and leave its proof to future research.
Appendix
A Technical definitions and results
Assumptions C3 and C4
The following definitions follow the exposition of Shapiro et al. (2021, Sec. 9.1.5). Let $\mathbb{T}\subset\mathbb{R}^{q}$ for some $q\in\mathbb{N}$. We firstly define
$\mathcal{T}_{\mathbb{T}}^{2}\left(\theta,\eta\right)=\left\{\tau\in\mathbb{R}^{q}:\text{dist}\left(\theta+t\eta+\frac{1}{2}t^{2}\tau,\mathbb{T}\right)=o\left(t^{2}\right),t\geq 0\right\}$
to be the second order tangent set of $\mathbb{T}$, at the point $\theta\in\mathbb{T}$, in the direction $\eta\in\mathcal{T}_{\mathbb{T}}\left(\theta\right)$. Here, $\mathcal{T}_{\mathbb{T}}\left(\theta\right)$ is the so-called the contingent cone to $\mathbb{T}$ at $\theta\in\mathbb{T}$, defined as the set of vectors $\eta\in\mathbb{R}^{q}$, such that there exist a sequences $\left(\eta_{j}\right)_{j\in\mathbb{N}}\subset\mathbb{R}^{q}$ and $\left(t_{j}\right)_{j\in\mathbb{N}}\subset\mathbb{R}_{\geq 0}$, where $\lim_{j\rightarrow\infty}\eta_{j}=\eta$ and $\theta+t_{j}\eta_{j}\in\mathbb{T}$, for each $j\in\mathbb{N}$. We then say that $\mathbb{T}$ is second order regular at $\theta^{*}\in\mathbb{T}$ if for any sequence $\left(\theta_{j}\right)_{j\in\mathbb{N}}\subset\mathbb{T}$ of the form
$\theta_{j}=\theta^{*}+t_{j}\eta+\frac{1}{2}t_{j}^{2}\tau_{j}\text{,}$
where $\lim_{j\rightarrow\infty}t_{j}=0$ and $\lim_{j\rightarrow\infty}t_{j}\tau_{j}=0$ for sequences $\left(t_{j}\right)_{j\in\mathbb{N}}\subset\mathbb{R}_{\geq 0}$ and $\left(\tau_{j}\right)_{j\in\mathbb{N}}\subset\mathbb{R}^{q}$, it follows that
$\lim_{j\rightarrow\infty}\text{dist}\left(\tau_{j},\mathcal{T}_{\mathbb{T}}^{2}\left(\theta^{*},\eta\right)\right)=0\text{.}$
Next, let $r:\mathbb{T}\rightarrow\mathbb{R}$ be a function of interest. We say that $r$ satisfies the quadratic growth condition at $\theta^{*}\in\mathbb{T}$ if there exists a constant $\mathfrak{C}>0$ and a neighborhood $\mathcal{N}\subset\mathbb{R}^{q}$ of $\theta^{*}$, such that
$r\left(\theta\right)\geq r\left(\theta^{*}\right)+\mathfrak{C}\left\|\theta-\theta^{*}\right\|^{2}\text{,}$
for all $\theta\in\mathcal{N}\cap\mathbb{T}$.
As a consequence of Bonnans and Shapiro (2000, Prop. 3.88), we have the fact that if
$\mathbb{T}=\left\{\theta\in\mathbb{R}^{q}:g_{u}\left(\theta\right)\leq 0\text{, }u\in\left[\mathfrak{m}\right],h_{v}\left(\theta\right)=0\text{, }v\in\left[\mathfrak{n}\right]\right\}$
for some $\mathfrak{m},\mathfrak{n}\in\mathbb{N}$, where each $g_{u},h_{v}:\mathbb{R}^{q}\rightarrow\mathbb{R}$ is twice continuously differentiable, and the point $\theta^{*}$ satisfies the so-called Mangasarian–Fromovitz constraint qualification, then $\mathbb{T}$ is second order regular at $\theta^{*}$. Here, the Mangasarian–Fromovitz constraint qualification can be stated as follows: the vectors $\left.\partial h_{v}\left(\theta\right)/\partial\theta\right|_{\theta=\theta^{*}}$ ($v\in\left[\mathfrak{n}\right]$) are linearly independent, and there exists a vector $\lambda\in\mathbb{R}^{q}$, such that $\lambda^{\top}\left.\partial h_{v}\left(\theta\right)/\partial\theta\right|_{\theta=\theta^{*}}=0$, for each $v\in\left[n\right]$, and $\lambda^{\top}\left.\partial g_{u}\left(\theta\right)/\partial\theta\right|_{\theta=\theta^{*}}<0$, for each $u\in\left[\mathfrak{m}\right]$, such that $g_{u}\left(\theta^{*}\right)=0$. In particular, $\theta^{*}\in\mathbb{T}$ is always second order regular if $\mathbb{T}$ is polyhedral, in the sense that $g_{u}$ and $h_{v}$ are affine functions of $\theta\in\mathbb{T}$, for each $u\in\left[\mathfrak{m}\right]$ and $v\in\left[\mathfrak{n}\right]$.
Using Shapiro et al. (2021, Props. 9.16 and 9.21), we have the facts that if $\theta^{*}\in\mathbb{T}$ is a global minimum of $r$, and $r$ is continuously differentiable at $\theta^{*}$, then $\eta^{\top}\left.\partial r\left(\theta\right)/\partial\theta\right|_{\theta=\theta^{*}}\geq 0$, for every $\eta\in\mathcal{T}_{\mathbb{T}}\left(\theta^{*}\right)$, and if $\mathbb{T}$ is second order regular and $r$ is twice continuously differentiable at $\theta^{*}$, then $r$ satisfies the quadratic growth condition at $\theta^{*}$, if and only if,
$\eta^{\top}\left.\frac{\partial^{2}r\left(\theta\right)}{\partial\theta\partial\theta^{\top}}\right|_{\theta=\theta^{*}}\eta+\inf_{\lambda\in\mathcal{T}_{\mathbb{T}}^{2}\left(\theta^{*},\eta\right)}\lambda^{\top}\left.\frac{\partial r\left(\theta\right)}{\partial\theta}\right|_{\theta=\theta^{*}}>0\text{,}$ (21)
for every $\eta\in\mathcal{C}\left(\theta^{*}\right)\backslash\left\{0\right\}$, where
$\mathcal{C}\left(\theta^{*}\right)=\left\{\eta\in\mathcal{T}_{\mathbb{T}}\left(\theta^{*}\right):\eta^{\top}\left.\frac{\partial r\left(\theta\right)}{\partial\theta}\right|_{\theta=\theta^{*}}=0\right\}$
is the so-called critical cone at the minimizer $\theta^{*}$. As noted in Shapiro (2000), the second term on the LHS of (21) is equal to zero if $\left.\partial r\left(\theta\right)/\partial\theta\right|_{\theta=\theta^{*}}=0$ or if $\mathbb{T}$ is polyhedral.
B Proofs
Proof of Theorem 1
Our proof largely follows the approach as that of Baudry (2015, Thm. 8.1). Firstly, the IID assumption together with A1 and A2 permits the application of a uniform strong law of large numbers (e.g., Shapiro et al., Thm. 9.60), which then implies that, for each $k$,
$\min_{\theta_{k}\in\mathbb{T}_{k}}R_{k,n}\left(\theta_{k}\right)\stackrel{{\scriptstyle[}}{{n}}\rightarrow\infty]{\text{a.s.}}{\longrightarrow}\min_{\theta_{k}\in\mathbb{T}_{k}}r_{k}\left(\theta_{k}\right)\text{,}$ (22)
by Shapiro et al. (2021, Prop. 5.2).
Let
$\mathcal{K}=\underset{k\in\left[m\right]}{\arg\min}\left\{\min_{\theta_{k}\in\mathbb{T}_{k}}r_{k}\left(\theta_{k}\right)\right\}$
and for each $k\notin\mathcal{K}$, let
$\epsilon=\frac{1}{2}\left\{\min_{\theta_{k}\in\mathbb{T}_{k}}r_{k}\left(\theta_{k}\right)-\min_{\theta_{k^{*}}\in\mathbb{T}_{k^{*}}}r_{k^{*}}\left(\theta_{k^{*}}\right)\right\}>0\text{.}$ (23)
Then, (22) and B1 imply that for every $\delta>0$,
$\left|\min_{\theta_{k}\in\mathbb{T}_{k}}R_{k,n}\left(\theta_{k}\right)-\min_{\theta_{k}\in\mathbb{T}_{k}}r_{k}\left(\theta_{k}\right)\right|\leq\frac{\epsilon}{3}\text{,}$ (24)
$\left|\min_{\theta_{k^{*}}\in\mathbb{T}_{k^{*}}}R_{k^{*},n}\left(\theta_{k^{*}}\right)-\min_{\theta_{k^{*}}\in\mathbb{T}_{k^{*}}}r_{k^{*}}\left(\theta_{k^{*}}\right)\right|\leq\frac{\epsilon}{3}\text{,}$ (25)
and $P_{k,n}\leq\epsilon/3$ all occur simultaneously with probability at least $1-\delta$, whenever $n$ is sufficiently large (i.e., $n\geq n_{\delta}$, for some large $n_{\delta}$). Thus,
$\min_{\theta_{k}\in\mathbb{T}_{k}}R_{k,n}\left(\theta_{k}\right)+P_{k,n}$ $\geq\min_{\theta_{k}\in\mathbb{T}_{k}}r_{k}\left(\theta_{k}\right)-\frac{\epsilon}{3}$
$=\min_{\theta_{k^{*}}\in\mathbb{T}_{k^{*}}}r_{k^{*}}\left(\theta_{k^{*}}\right)+\frac{5\epsilon}{3}$
$\geq\min_{\theta_{k^{*}}\in\mathbb{T}_{k^{*}}}R_{k^{*},n}\left(\theta_{k^{*}}\right)+P_{k^{*},n}+\epsilon$
also occurs with probability at least $1-\delta$, for sufficiently large $n$. But since $\epsilon>0$, we have the fact that
$\min_{\theta_{k}\in\mathbb{T}_{k}}R_{k,n}\left(\theta_{k}\right)+P_{k,n}>\min_{\theta_{k^{*}}\in\mathbb{T}_{k^{*}}}R_{k^{*},n}\left(\theta_{k^{*}}\right)+P_{k^{*},n}\text{,}$ (26)
and thus by definition of (2), $\hat{K}_{n}\neq k$ occurs with probability at least $1-\delta$, for large $n$, for each $k\notin\mathcal{K}$. Because there is only a finite number of $k\notin\mathcal{K}$, we have the fact that
$\lim_{n\rightarrow\infty}\text{Pr}\left(\hat{K}_{n}\notin\mathcal{K}\right)=0\text{.}$
Next, we consider when $k\in\mathcal{K}$ but $k>k^{*}$. From A1 and A2, Lemma 3 implies that for each $k\in\mathcal{K}$,
$\sqrt{n}\left(\min_{\theta_{k}\in\mathbb{T}_{k}}R_{k,n}\left(\theta_{k}\right)-\min_{\theta_{k}\in\mathbb{T}_{k}}r_{k}\left(\theta_{k}\right)\right)=O_{\text{Pr}}\left(1\right)$
which implies that
$\sqrt{n}\left(\min_{\theta_{k}\in\mathbb{T}_{k}}R_{k,n}\left(\theta_{k}\right)-\min_{\theta_{k^{*}}\in\mathbb{T}_{k^{*}}}R_{k^{*},n}\left(\theta_{k^{*}}\right)\right)=O_{\text{Pr}}\left(1\right)\text{,}$
since $\min_{\theta_{k^{*}}\in\mathbb{T}_{k^{*}}}r_{k^{*}}\left(\theta_{k^{*}}\right)=\min_{\theta_{k}\in\mathbb{T}_{k}}r_{k}\left(\theta_{k}\right)$, by definition of $\mathcal{K}$. Thus, for each $k$ and $\delta>0$, there exists a constant $\mathfrak{M}>0$, such that
$\sqrt{n}\left|\min_{\theta_{k}\in\mathbb{T}_{k}}R_{k,n}\left(\theta_{k}\right)-\min_{\theta_{k^{*}}\in\mathbb{T}_{k^{*}}}R_{k^{*},n}\left(\theta_{k^{*}}\right)\right|\leq\mathfrak{M}$
with probability at least $1-\delta$, for sufficiently large $n$. On the other hand, by B2, for each $\delta>0$ and $\mathfrak{M}>0$,
$\sqrt{n}\left\{P_{k,n}-P_{k^{*},n}\right\}>\mathfrak{M}$
with probability at least $1-\delta$, for sufficiently large $n$. Thus, for sufficiently large $n$, with probability at least $1-2\delta$, we have
$\min_{\theta_{k^{*}}\in\mathbb{T}_{k^{*}}}R_{k^{*},n}\left(\theta_{k^{*}}\right)-\min_{\theta_{k}\in\mathbb{T}_{k}}R_{k,n}\left(\theta_{k}\right)\leq\frac{\mathfrak{M}}{\sqrt{n}}<P_{k,n}-P_{k^{*},n}$
and thus
$\min_{\theta_{k^{*}}\in\mathbb{T}_{k^{*}}}R_{k^{*},n}\left(\theta_{k^{*}}\right)+P_{k^{*},n}<\min_{\theta_{k}\in\mathbb{T}_{k}}R_{k,n}\left(\theta_{k}\right)+P_{k,n}\text{,}$
for each $k\in\mathcal{K}$, such that $k>k^{*}$. Again, since there is a finite number of elements of $\mathcal{K}$, and by definition of (2), we have
$\lim_{n\rightarrow\infty}\text{Pr}\left(\hat{K}_{n}\in\mathcal{K},\hat{K}_{n}>k^{*}\right)=0.$
Finally, we obtain the desired result via the limit:
$\lim_{n\rightarrow\infty}\text{Pr}\left(\hat{K}_{n}\neq k^{*}\right)=\lim_{n\rightarrow\infty}\left\{\text{Pr}\left(\hat{K}_{n}\notin\mathcal{K}\right)+\text{Pr}\left(\hat{K}_{n}\in\mathcal{K},\hat{K}_{n}>k^{*}\right)\right\}=0\text{.}$
Proof of Theorem 2
Under A1, A2, and C1–C5, Shapiro (2000, Thm. 4.4) implies the following result (via an application of Shapiro et al., 2021, Thm. 9.56).
Lemma 3.
Assume that $\left(X_{i}\right)_{i\in\left[n\right]}$ is an IID sequence, and that A1, A2, and C1–C5 hold, for each $k\in\left[m\right]$. Then, for each $k$,
$n\left(\min_{\theta_{k}\in\mathbb{T}_{k}}R_{k,n}\left(\theta_{k}\right)-R_{k,n}\left(\theta_{k}^{*}\right)\right)\rightsquigarrow\frac{1}{2}\varphi_{k}\left(Z_{k}\right)\text{,}$ (27)
where $Z_{k}$ is normally distributed with mean zero and covariance matrix
$\mathrm{E}\left\{\left[\left.\frac{\partial\ell_{k}\left(X;\theta_{k}\right)}{\partial\theta_{k}}\right|_{\theta_{k}=\theta_{k}^{*}}\right]\left[\left.\frac{\partial\ell_{k}\left(X;\theta_{k}\right)}{\partial\theta_{k}}\right|_{\theta_{k}=\theta_{k}^{*}}\right]^{\top}\right\}-\left[\left.\frac{\partial r_{k}\left(\theta\right)}{\partial\theta_{k}}\right|_{\theta_{k}=\theta_{k}^{*}}\right]\left[\left.\frac{\partial r_{k}\left(\theta\right)}{\partial\theta_{k}}\right|_{\theta_{k}=\theta_{k}^{*}}\right]^{\top}$
and
$\varphi_{k}\left(\zeta\right)=\inf_{\eta\in\mathcal{C}\left(\theta^{*}\right)}\left\{2\eta^{\top}\zeta+\eta^{\top}\left.\frac{\partial^{2}r\left(\theta\right)}{\partial\theta_{k}\partial\theta_{k}^{\top}}\right|_{\theta_{k}=\theta_{k}^{*}}\eta+\inf_{\lambda\in\mathcal{T}_{\mathbb{T}_{k}}^{2}\left(\theta_{k}^{*},\eta\right)}\lambda^{\top}\left.\frac{\partial r_{k}\left(\theta\right)}{\partial\theta_{k}}\right|_{\theta_{k}=\theta_{k}^{*}}\right\}\text{.}$
As per the use of Lemma 1, we only require Lemma 3 to obtain the fact that
$n\left(\min_{\theta_{k}\in\mathbb{T}_{k}}R_{k,n}\left(\theta_{k}\right)-R_{k,n}\left(\theta_{k}^{*}\right)\right)=O_{\text{Pr}}\left(1\right)\text{,}$
for each $k\in\left[m\right]$. Recalling the definition of $\mathcal{K}$ from the previous proof, via C5 and Lemma 3, we then have
$n\left(\min_{\theta_{k}\in\mathbb{T}_{k}}R_{k,n}\left(\theta_{k}\right)-\min_{\theta_{k^{*}}\in\mathbb{T}_{k^{*}}}R_{k^{*},n}\left(\theta_{k^{*}}\right)\right)$
$=$ $n\left(\min_{\theta_{k}\in\mathbb{T}_{k}}R_{k,n}\left(\theta_{k}\right)-R_{k,n}\left(\theta_{k}^{*}\right)\right)-n\left(\min_{\theta_{k^{*}}\in\mathbb{T}_{k^{*}}}R_{k^{*},n}\left(\theta_{k^{*}}\right)-R_{k^{*},n}\left(\theta_{k^{*}}^{*}\right)\right)$
$-n\left(R_{k,n}\left(\theta_{k}^{*}\right)-R_{k^{*},n}\left(\theta_{k^{*}}^{*}\right)\right)$
$=$ $O_{\text{Pr}}\left(1\right)+O_{\text{Pr}}\left(1\right)+O_{\text{Pr}}\left(1\right)=O_{\text{Pr}}\left(1\right)\text{,}$
for each $k\in\mathcal{K}$. The remainder of the proof follows in the same way as that of Theorem 1.
Proof of Theorem 3
The proof follows in the same manner as that of Theorem 1. By (22) and B1, for each $k\notin\mathcal{K}$, we have both (24) and (25), almost surely when $n$ is sufficiently large, for $\epsilon$ defined by (23). This then implies (26), almost surely, for each $k\notin\mathcal{K}$, and thus, we can conclude that
$\text{Pr}\left(\lim_{n\rightarrow\infty}\hat{K}_{n}\notin\mathcal{K}\right)=0\text{,}$
since $\left[m\right]$ is finite.
Next, when A1 and A2 hold, and when $k\in\mathcal{K}$ and $k>k^{*}$, Lemma 2 implies that
$\sqrt{\frac{n}{\log_{+}^{\left(2\right)}\left(n\right)}}\left(\min_{\theta_{k}\in\mathbb{T}_{k}}R_{k,n}\left(\theta_{k}\right)-\min_{\theta_{k}\in\mathbb{T}_{k}}r_{k}\left(\theta_{k}\right)\right)=O_{\text{a.s.}}\left(1\right)$
and hence
$\sqrt{\frac{n}{\log_{+}^{\left(2\right)}\left(n\right)}}\left(\min_{\theta_{k}\in\mathbb{T}_{k}}R_{k,n}\left(\theta_{k}\right)-\min_{\theta_{k^{*}}\in\mathbb{T}_{k^{*}}}R_{k^{*},n}\left(\theta_{k^{*}}\right)\right)=O_{\text{a.s.}}\left(1\right)\text{,}$
by definition of $\mathcal{K}$. Thus, for each $k\in\mathcal{K}$, there exists a constant $\mathfrak{M}>0$ such that
$\sqrt{\frac{n}{\log_{+}^{\left(2\right)}\left(n\right)}}\left|\min_{\theta_{k}\in\mathbb{T}_{k}}R_{k,n}\left(\theta_{k}\right)-\min_{\theta_{k^{*}}\in\mathbb{T}_{k^{*}}}R_{k^{*},n}\left(\theta_{k^{*}}\right)\right|\leq\mathfrak{M}$ (28)
for sufficiently large $n$, almost surely. Then, by B2∗∗, for each $\mathfrak{M}>0$
$\sqrt{\frac{n}{\log_{+}^{\left(2\right)}\left(n\right)}}\left\{P_{k,n}-P_{k^{*},n}\right\}>\mathfrak{M}$
for sufficiently large $n$, almost surely, which together with (28), implies
$\min_{\theta_{k^{*}}\in\mathbb{T}_{k^{*}}}R_{k^{*},n}\left(\theta_{k^{*}}\right)-\min_{\theta_{k}\in\mathbb{T}_{k}}R_{k,n}\left(\theta_{k}\right)\leq\sqrt{\frac{\log_{+}^{\left(2\right)}\left(n\right)}{n}}\mathfrak{M}<P_{k,n}-P_{k^{*},n}$
and hence
$\min_{\theta_{k^{*}}\in\mathbb{T}_{k^{*}}}R_{k^{*},n}\left(\theta_{k^{*}}\right)+P_{k^{*},n}<\min_{\theta_{k}\in\mathbb{T}_{k}}R_{k,n}\left(\theta_{k}\right)+P_{k,n}\text{,}$
for sufficiently large $n$, almost surely, as required. Again, since $\mathcal{K}$ is finite, the intersect of the required almost sure events is also almost sure, and hence
$\text{Pr}\left(\lim_{n\rightarrow\infty}\hat{K_{n}}\in\mathcal{K},\lim_{n\rightarrow\infty}K_{n}>k^{*}\right)=0$
which then implies the desired result
$\text{Pr}\left(\lim_{n\rightarrow\infty}\hat{K}_{n}\neq k^{*}\right)=\text{Pr}\left(\lim_{n\rightarrow\infty}\hat{K}_{n}\notin\mathcal{K}\right)+\text{Pr}\left(\lim_{n\rightarrow\infty}\hat{K_{n}}\in\mathcal{K},\lim_{n\rightarrow\infty}K_{n}>k^{*}\right)=0\text{.}$